3) A wireless receiver needs to meet the sensitivity requirement of -90 dBm. The channel bandwidth is 1 MHz. The data rate is 0.5 Mbps. The antenna temperature is 290 K. (Please show all steps to receive full credits) a) If Eb/No requirement is 10 dB. What is the SNR requirement in dB? (10 pts) b) [independent of (a)] If SNR requirement is 10 dB, what is the NF requirement in dB? (15 pts)
Answer #1

  Eb fb No B We have Smin - kToB(NF) and We have SNR-.

\mathrm{where\;\;\;S_{min}=Receiver\;sensitivity}

  \mathrm{\left ( \frac{S}{N} \right )_{min}=Minimum\;Signal\;to\;Noise\;ratio(SNR)}

\mathrm{k=Boltzmann's\;constant=1.38\times10^{-23}Joule/^{0}K}

\mathrm{T_0=Absolute\; temperature\;in\;^{0}K }

  \mathrm{B=Bandwidth}

\mathrm{NF=Noise\;figure}

f-Data rate

\mathrm{Here\;given\;\;S_{min}=-90\;dBm}\\

  \mathrm{B=1\;MHz}

  \mathrm{T_0=290\;K}

  \mathrm{Data\;rate=0.5\;Mbps}

\mathrm{\mathbf{(a)}}

D. D No B We have SNR -

Where fb  is data rate and it was given as  fb = 0.5 Mbps

\mathrm{Given\;\;\frac{E_b}{N_0}=10\; dB=10}

  \therefore \mathrm{SNR=(10)\left (\frac{0.5}{1} \right )}

= 5

= 10 log(5)

SNR = 6.989 dB

\mathrm{\mathbf{(b)}}

  10 Given SNR= = 10 dB

  \therefore \mathrm{NF=\frac{S_{min}}{(SNR)kT_0B}}

Smin = -90 dBm = 10-12 Watt

SNR = 10 dB = 10

10-12 (10)(1.38 × 10-23)(290) (1 × 106) NF =

\mathrm{=24.9875}

10 log(24.9875) 13.977 dB

\therefore \mathrm{NF=13.977\;dB}